OpenStudy (anonymous):

solve 1/n^2+1/n=1/2n^2 please help me get through this and show work so i can get a good understanding

OpenStudy (turingtest):

\[\frac1{n^2}+\frac1n=\frac1{2n^2}\]did I type it right?

OpenStudy (turingtest):

@sassyjazz could you please tell my if I have typed the problem correctly? it's hard to help you if you don't stick around

OpenStudy (anonymous):

Multiply through to eliminate denominator. Since they are all powers/factors of n, find the largest value and multiply across (2n^2). Move all terms to one side. Solve the quadratic equation for n.

OpenStudy (anonymous):

YES UU DID

OpenStudy (anonymous):

SORRY COMPUTER WAS MESSING UP

OpenStudy (anonymous):

\[1/n ^{2}+1/n=1/(2*n ^{2})\] multiply by n^2 1+n=0.5 n=-0.5

OpenStudy (anonymous):

bingo

OpenStudy (turingtest):

@sassyjazz you still there? do you understand any better now?

OpenStudy (anonymous):

KINDA

OpenStudy (turingtest):

the idea is to find the GCF of all the denominators once we find that, if we multiply both sides by the GCF all the fractions will disappear

OpenStudy (turingtest):

so what are the denominators?\[n^2,~n,~\text{ and } 2n^2\]do you know how to find the greatest common factor of those numbers?

OpenStudy (anonymous):

WOULD IT B ^2